Having Problems solving this solution
The Everton College store paid $1626 for a order of 45 calculators. The store paid $10 for each scientific calculator. The others, all graphing calculators, cost the store $52 each. How many of each type of calculator was ordered?
The store ordered___? scientific calculators and ____graphing calculators?

Answer by mananth(16949)

$10 for each scientific calculator. ------------ x numbers
graphing calculators, cost $52 each.------------ y numbers
..
total = 45 calculators
...
value = $1626
...
x+y =45
10x+52y=1626
...
multiply first equation by -10
-10x-10y=-450
add to second equation
52y-10y=1626-450
42y=1176
/42
y = 1176/42
y= 28 graphing calculators
balance 17 = scientific calculators
...
m.ananth@hotmail.ca